2. [5 pts] You hear a loud shout from your friend at 80 db. If you move twice as far away from your friend, how loud is the shout?
Answer #1

We will use the sound attenuation formula

Where

S1 = Sound pressure level at point 1 = 80 db

r2 = distance of point 2 = 2 x distance at point 1 = 2r1

So, Sound pressure level at point 2, S2 = 80 - 20 log(2r1/r1) db = 80 -  20 log(2) db = 80 - 6.02 db = 73.98 db

So, the shout will now be heard at 73.98 db-----------------------(Answer)